OpenStudy (anonymous):

Find a third degree polynomial that has zeros at x=-1,1, and 2. whose x term has coefficient 3. help!? Ill give out medals!

OpenStudy (cwrw238):

it will be of the form a(x+1)(x - 1)(x - 2) where a is a constant

OpenStudy (anonymous):

really? because for like a 4th degree one where x=0,1,-1, and -2 and the x^2 term has coefficient of 5. the answer is : -5x^4-10x^3+5x^2+10x.. Im pretty confused..

OpenStudy (anonymous):

oh is that the form and then i just put the constant in front and it will get an answer?

OpenStudy (cwrw238):

expanding we get a(x + 1)(x^2 - 3x + 2) = a( x^3 - 3x^2 + 2x + x^2 - 3x + 2) = a(x^3 -2x^2 - x + 2) this should help

OpenStudy (anonymous):

... ugh i know im probably being stupid or something, but im on no sleep. So just asking.. what is this? Is that the answer? an equation i use to get the answer?..

OpenStudy (anonymous):

@cwrw238 ?

OpenStudy (cwrw238):

sorry - I was called away

OpenStudy (anonymous):

oh.

OpenStudy (cwrw238):

right so the coefficient of x must be 3 - that makes a = 3 so you have to multiply each term in the brackets by 3 and you have your polymimial

OpenStudy (cwrw238):

* polynomial lol

OpenStudy (anonymous):

oh, ok. lol. Thanks a lot!. can i ask you 1 more question?

OpenStudy (cwrw238):

ok

OpenStudy (anonymous):

so i have to find the possible rational zeros for x^4-4x^3+3x^2+4x-4. I think those are +-1, 2, and 4. Then i have to factor it. So i guess i just try them out to see which ones work and then factor?

OpenStudy (cwrw238):

they look right - because the last term is -4 yes just check them out first

OpenStudy (anonymous):

ok. but question.. for another problem earlier which ended in -3 and the first erm was 2. The zero's were ony 1,-1/2, and -3. Not +- 1,3, -3/2, or /1,2. Why did some just dissapear. And then the polynomial ends up being factored by x-1 which i thought should be x+1?

OpenStudy (cwrw238):

not all the factors are necessarily zeros if one zero is 1 then it will be divisible by x - 1 (that's the Factor Theorem)

OpenStudy (anonymous):

oh. so if it was -1 it would be x+1?

OpenStudy (cwrw238):

yes

OpenStudy (cwrw238):

I checked out -1,1 and 2 for your other equation and they fit OK

OpenStudy (anonymous):

yea, i was checking them now actually. Was on 4. lol

OpenStudy (anonymous):

I literally have this class in about 20 minutes so If you could help with 1 more question...I got another one. like the one we did before though my signs were messed up. But what would be the equation for this: 5th degree. zero of multiplicity 2 at x=1, zero at x=4 and factor of x^2+x+1? that one just i have no idea.

OpenStudy (cwrw238):

4 does not fit -according to me

OpenStudy (cwrw238):

maybe 2 is a repeat root in which case the factors are (x+1)(x-1)(x-2)^2

OpenStudy (cwrw238):

yes 2 is a repeated root

OpenStudy (anonymous):

so for the 5th degree one. that is the equation? and i just solve that?

OpenStudy (anonymous):

oh yea. I see that.

OpenStudy (anonymous):

wow. ok. I think i actually get this stuff now. :)

OpenStudy (cwrw238):

is best stated as multiplicity 2 not repeated as I did

OpenStudy (anonymous):

yea. I am squaring the final answer though right? after i distribute each one..

OpenStudy (cwrw238):

yes

OpenStudy (cwrw238):

gotta go right now sorry

OpenStudy (anonymous):

its ok. you helped alot! thanks so much!! Yea, and i got x^3-2x^2-x+2 and then square that and i should be set. Thanks again!